OpenStudy (anonymous):

3) Write the following expression as the sine, cosine, or tangent of an angle: cos45°cos120° - sin45°sin120 solve in steps

OpenStudy (anonymous):

cos(a+b) = cos a.cosb - sin a. sinb so the given expression = cos (45 - 120) = cos(-75) =cos75